OpenStudy (chycora):

At a baseball game, the probability that a fan brings a glove is 0.08. The probability that a fan is left-handed is 0.12. What is the probability that a randomly selected fan has a baseball glove and is left-handed? Round your answer to the nearest hundredth, if necessary. A. 0.67 B. 0.04 C. 0.01 D. 0.20

OpenStudy (amistre64):

what are your thoughts?

OpenStudy (chycora):

maybe a

OpenStudy (amistre64):

more to the point, why do you think its A? whats your process

OpenStudy (chycora):

idk if it's exactly a i just guessed but i gt something else see i mulitiplied 0.12* 0.8 and gt 0.096 u see and if i round that to the nearest hundredths it will be 0.1 so yea

OpenStudy (chycora):

do u see @amistre64 i'm so confused

OpenStudy (amistre64):

.08 * .12 is an AND construction i agree P(A|B) = P(AnB)/P(B) P(B) * P(A|B) = P(AnB) if A and B are independant, then P(A|B) = P(A)

OpenStudy (amistre64):

are the events independant?

OpenStudy (chycora):

no i think

OpenStudy (chycora):

am i correct @amistre64

OpenStudy (chycora):

i think i kn the answer is it C

OpenStudy (amistre64):

i dont kow what the right answer is, this is all we have for the information.